| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 1992 001 00 04 05 Motor #10: appears stuck with a reference sensor activated. | 2 | 1 | B21924.192 | TORON | #192 | 05:25:52 | 1 | -1 | |
| hg: | Lamp not on... hg test terminated F1= 0 F4= 10 | 2 | 1 | B21924.192 | TORON | #192 | 06:43:15 | 1 | -1 | |
| hg: | Lamp not detected f1= 0 f4= 10 | 2 | 1 | B21924.192 | TORON | #192 | 06:43:16 | 1 | -1 | |
| hp: | Spectrometer alignment trouble. | 1 | 1 | B21924.192 | TORON | #192 | 19:38:07 | 1 | -1 |
| Serial | YYYYJJJ | HG large | max correction |
|---|---|---|---|
| 192 | 2024219 | 33 | 6.4 |
| Serial | YYYYJJJ | HP large | max correction |
|---|---|---|---|
| 192 | 2024219 | 13 | 80.0 |
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
|---|---|---|---|---|---|---|---|---|---|---|
| rl: | 1992 001 00 04 04 Reset Warm -> Tepid from 1992 001 00 39 36 until 1992 001 00 01 27 . Counts: 1,0 | 1 | 3 | B21924.192 | TORON | #192 | 05:25:50 | 1 | -1 | |
| rl: | 1992 001 00 04 05 Motor #10: appears stuck with a reference sensor activated. | 1 | 3 | B21924.192 | TORON | #192 | 05:25:52 | 1 | -1 | |
| rl: | 1992 001 00 04 05 Motor #11: appears stuck with a reference sensor activated. | 1 | 3 | B21924.192 | TORON | #192 | 05:25:55 | 1 | -1 | |
| rl: | 1992 001 00 05 01 New value set for watermark SUPPLY.VOLTAGE.HIGH | 3 | 3 | B21924.192 | TORON | #192 | 05:25:58 | 1 | -1 | |
| dp: | Doing dispersion with internal HG. | 1 | 99 | B21924.192 | TORON | #192 | 09:14:37 | 1 | -1 | |
| dt: | HOME key pressed | 2 | 99 | B21924.192 | TORON | #192 | 11:40:46 | 1 | -1 | |
| pd: | Going to menu | 2 | 99 | B21924.192 | TORON | #192 | 11:41:02 | 1 | -1 | |
| re: | Azimuth discrepancy = -16000 | 1 | 99 | B21924.192 | TORON | #192 | 18:16:17 | -16000 | -30 | 30 |
| hg: | DEL key pressed | 2 | 99 | B21924.192 | TORON | #192 | 19:56:11 | 1 | -1 |
| _jday | _loc | _serial | _skc | dt | ed | hg | hp | rs | sl | uv |
|---|---|---|---|---|---|---|---|---|---|---|
| 219 | TORONTO | #192 | inside,menu,Must do HG now,Waiting until ZA = 170 now 118.2,Waiting until ZA = 170 now 118.7,Waiting until ZA = 170 now 119.1,Waiting until ZA = 170 now 119.4,Waiting until ZA = 170 now 119.7,Waiting until ZA = 170 now 119.8,Waiting until ZA = 170 now 119.9,Waiting until ZA = -100 now -100.1,Waiting until ZA = -100 now -101.6,Waiting until ZA = -100 now -103.1,Waiting until ZA = -80 now -81,Waiting until ZA = -85 now -85.5,Waiting until ZA = -85 now -87.2,Waiting until ZA = -85 now -88.7,Waiting until ZA = -85 now -90.8,Waiting until ZA = -85 now -92.4,Waiting until ZA = -85 now -94 | 11 | 2 | 37 | 41 | 10 | 10 | 1 |
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
|---|---|---|---|---|
| 192 | 33.8 | 7.8 |